i do not wish for the answer i just need help working it out
the problem is -7/16 - 3/8

Found 2 solutions by stanbon, ilana:
Answer by stanbon(48569)

When you subtract or add fractions you need to
have the same denominator in the fractions.
So you determine the "least common denominator"
be finding the smallest number into which the
denominators you have can divide evenly. That
number in your case is 16.
The 1st fraction already has that denominator so
you leave it alone.
For the 2nd fraction to have a denominator of 16
you would have to multiply it by 2. So you multiply
both top and bottom of that fraction by 2 to get
the following:
-7/16 - 6/16
Now you subtract the numerators and write the result
over 16 to get:
-13/16.
Hope this helps.
Cheers,
Stan H.

Answer by ilana(305)

That's a great strategy- not wanting the answer I mean.
Ok, so whenever we want to add or subtract fractions, we need to start with a common denominator. Sometimes this is already present (like 4/5 - 2/5), and sometimes we need to get our own (like 4/5 - 2/15). If we have a common denominator, we simply combine the numerators (4/5 - 2/5 = (4-2)/5 = 2/5). If we need to find our own, we have more cases.
The first is that one denominator is a factor of the other. (5 is a factor of 15 since 3*5=15. So for 4/5 - 2/15 we multiply 4/5 by 3/3, which we can do since 3/3 = 1, and get 12/15 - 2/15 =10/15 =2/3, our answer here).
Another possibility is that both denominators have a common factor. (If we have 2/10 + 4/15, we see 10=2*5 and 15=3*5, so we can get a common denominator of 2*3*5=30. So we multiply 2/10 by 3/3 and multiply 4/15 by 2/2 and get 6/30 + 8/30 = 14/30 = 7/15).
Finally, it is possible that the denominators have no common factors. (If we have -2/5 - 1/12, we see 5 and 12 have no common factors, so we can get a common denominator by multiplying 5 and 12 to get 60. So, we multiply -2/5 by 12/12 and -1/12 by 5/5 and get (-24/60)+(-5/60)= -29/60).
*Note: IN this last case, we changed -2/5 - 1/12 to -2/5 + -1/12.
**Note: This last strategy of multiplying the denominators works in all cases, but in the previous cases it will give us unnecessarily large numbers to work with (With 2/10+4/15, we would get (2*15/10*15)+(4*10/15*10) = (30+40)/150 = 70/150 = 7/15).